Test

For a closed economy, GDP is $13 trillion, consumption is $9 trillion, taxes are $4 trillion and the government runs a surplus of $1 trillion. What is national saving?

a.$1 trillion

b.$5 trillion

c.$2 trillion

d.$3 trillion



In the economy of Russia in 2020, consumption was $3500, exports were $200, GDP was $7200, government purchases were $2500, and investment was $2,000. What were Russia’s imports in 2020?

a.$14,400

b.$800

c.$8200

d.$1,000



Suppose that in a closed economy GDP is equal to 10,000, taxes are equal to 3,500 consumption equals 6,000 and government purchases equal 2,000. What are private saving, public saving, and national saving?

a.2,500, 1,000, and 1,500, respectively

b.2,000, 2,000, and 1,500, respectively

c.2,000, 500, and 1,500, respectively
Answer #1

1.

The answer is Option A.
National saving is a portion of GDP rather than spent in an economy. It is calculated by dividing the difference between a nation's GDP and income divided by income. Here we have GDP = 13 trillion and Consumption spending= 9 trillion. Therefore private saving = 13-9=4. If we deduct tax and add government surplus, we get national savings. Ie, 4-4= 0, 0-1=1 trillion.


2.

GDP = Consumption + govt purchases + investment + net exports ( export - imports )

7200 unit = 3500+2500+2000+ (200-imports )

7200 - 8000 = 200- imports

-800- 200 = - imports

So Imports is equal to 1000 unit.


3.


answered by: Allen